How do I solve the following system of linear equation
5x + 4y = 40
x + 2y = 11

Answer by Alan3354(69443)

How do I solve the following system of linear equation
5x + 4y = 40
x + 2y = 11
-----------
There's more than one way. I like determinants (I have an Excel sheet that does them).
5 4 -40
1 2 -11
Det = (5*2)-(1*4) = 6
x*Det = 4*(-11)-2*(-40) = 36
x = 6
-----------
y*Det = -(5*(-11) - 1*(-40)) = 15
y = 2.5